The Model F extender fits with the 2 inch heavy duty hitches.

The contract calls for 201 Model F extenders per week to be delivered in equal installments over the 16 weeks of the contract. The goal of Alpha Assemblies is to work 40 hours per week. The actual work time for completing the Model F extenders has been broken down by process in the table below. Also provided is the anticipated learning rate for each process. All processes must be performed in sequence and each step has its own separate and unique workcenter. To achieve the goal of working 40 hours per week or less, the cycle time must be lower than the takt time.

Process Time Required per Unit Predecessor Task Learning Rate
A 10 88
B 11 A 87
C 18 B 86
D 9 C 82
E 15 D 84
F 8 E 81
G 9 F 84
H 16 G 85
I 10 H 86

Part of the answer:

Process

Time Required per Unit Predecessor Task Learning Rate Minutes(week2) Minutes(week4) Minutes(week8) Minutes(week16)
A 10 88% 8.8 7.744 6.81472 5.996954
B 11 A 87% 9.57 8.3259 7.243533 6.301874
C 18 B 86% 15.48 13.3128 11.449008 9.846147
D 9 C 82% 7.38 6.0516 4.962312 4.069096
E 15 D 84% 12.6 10.584 8.89056 7.46807
F 8 E 81% 6.48 5.2488 4.251528 3.443738
G 9 F 84% 7.56 6.3504 5.334336 4.480842
H 16 G 85% 13.6 11.56 9.826 8.3521
I 10 H 86% 8.6 7.396 6.36056 5.470082

For week 16 how much better (lower so answer will be positive) or worse (higher so answer will be negative) is the cycle time for Model F than the takt time? Note: learning is applied to the batch quantity per week. Do not try to break out the units within a week.

Hint: The answer is 2.0942
Answer #1

Cycle Time = Bottleneck of Activity for Week16

Cycle Time = 9.846147 minute (since it has the largest activity time)

Demand Per Week = 201

Available Time per week = 40*6 = 2400 minutes

Takt Time = Available Time Per Week/Demand PEr Week

Takt Time = 2400/201

Takt Time = 11.9403

Cycle Time lower than the Takt Time

Difference = 11.9403 - 9.846147

Difference = 2.094152

Difference = 2.0942 minutes